OpenStudy (precal):

TI-89 question I am using my ti-89 to find a trig derivative, sin^2(x) it is giving me (pi/90)sin(x)cos(x)

ganeshie8 (ganeshie8):

you must be in degrees mode, the calculator is converting degrees to radians for taking derivative it seems

ganeshie8 (ganeshie8):

maybe change it to radian mode and try ?

OpenStudy (precal):

yes you are correct. I also, had to make sure it was in exact mode.

ganeshie8 (ganeshie8):

\(\large (\sin (x~\text{degrees}))' = (\sin (x\times \dfrac{\pi}{180} ~\text{radians}))' = \dfrac{\pi}{180} \cos(x~\text{degrees})\)